# Problem Solving Through Human-AI Preference-Based Cooperation

### v3

@@ 4 Challenges and practical issues
Given that human experts are costly and have limited time, LLM-based simulation of human-AI interaction may facilitate large-scale evaluation (Tamoyan et al., 2024). Even though our four evaluation criteria seem to demand human evaluation, we conjecture that the development of artificial critic models (McAleese et al., 2024), with human value alignment, will be an important research direction in the future.
Figure 3: Instantiation of HAI-Co2 for the problem of building a double pendulum simulation. A solution is co-constructed through human-AI cooperation as follows. The interaction between the human and the AI is shown in the red box on the right, top to bottom. We define the co-construction space on three levels of hierarchy: Specification space, UML space, and Python program space. The user starts by specifying the (potentially incomplete) problem to solve. The Specification module (green rectangle) generates a pair of candidate specifications of the software to build (S1 and S2) and presents them to the user. The user expresses their preference in two manners: i) they choose one of the candidate solutions (S1) as better than the other, and ii) provide partial edits to the specification directly. The Preference decoder module (purple rectangle) extracts preference rules from the interaction context. Based on the decoded preferences, the Specification module generates a new pair of candidate specifications (only one shown for space reasons), from which the user chooses one. The UML generation module serves as a generator of refinements from specification space to UML space and generates a set of four UMLs from the specification selected. The Preference-based search module then runs a tournament-based search among these UML candidates: a pair of UMLs are compared against the specification and the decoded user preferences and one is chosen. Two finalist UMLs from the tournament are then used by the Code generation module (pink rectangle) to generate two candidate Python programs. These programs are presented to the user again for their feedback.
5 An exemplary simulation of HAI-Co2
@@
In this section, we provide a case study of HAI-Co2, tailored to code generation as a co-construction problem. This case study does not claim scientific rigor on its own; instead, we use it to demonstrate what prior findings (see Section 1 and Section 2) already establish. We do not provide a complete implementation of HAI-Co2; in particular, the following are not included in the case study: neighborhood structure of the solution space, preference extraction from actual human participation, dedicated utility function tailored toward the expert problem. Instead, we emulate the intended behavior of a complete implementation using prompted LLMs, with the goal of motivating the practicality of HAI-Co2.
@@
To facilitate the exploration of the candidate solution space, we generate multiple solution representations on different abstraction levels by setting a high decoding temperature in the respective generation modules and sampling multiple responses. Intuitively, we seek to exploit earlier findings that a highly stochastic generation regime facilitates better novelty (Wang et al., 2023a). Furthermore, this imposed stochasticity can be interpreted as the notion of neighborhood in the respective spaces: one can treat two solutions sampled from the same input context of the generation module as neighbors; distance between two different input contexts can be measured by edit distance. Although we do not explicitly specify such geometric structure in this example, the search strategy uses it implicitly.
@@
The preference-based search among the candidate UMLs is implemented as a tournament by iteratively declaring one among a pair as the winner of a round. After a logarithmic order of such rounds ($\log_{2}n$ being the depth of the tournament tree for $n$ candidate UMLs), the Preference-based search module comes up with a final pair and a summary of preference justifications. [^6] See https://subha0009.github.io/ExAIC-Interactions-10A2/PreferenceLayer.html for the tournament on the candidate UMLs generated Note that although we seek to minimize the cost of human intervention in this step by automating preference-based ranking, one can envision the human expert providing their judgement on these UMLs. In such a setup, the Preference-decoding module can be used to explicitly adapt to such gold preference examples.
@@
We recruited a total of 14 participants for this evaluation; each is a doctoral student in Natural Language Processing, so that expert preferences in code generation can be understood. On all five criteria, the majority of evaluators rate HAI-Co2 higher than GPT-4 Turbo. 12 (85.7%) evaluators find that HAI-Co2 better captures the initial preferences of the user (Q1). 11 (78.6%) agree that it can adapt better to preference switching (Q2). 11 (78.6%) see HAI-Co2 as superior on precision (Q3) and completeness of modification (Q4). 12 (85.7%) suggest HAI-Co2 is better suited for software development (Q5). We present detailed justifications for the responses provided by the evaluators in Appendix B.
@@
Limitations and further improvements. The immediate improvements we observe are prevalent without any dedicated implementation – neither of the two refinement maps (from natural language to UML and from UML to Python) nor of the preference-based search policy. In this implementation, we do not equip the co-construction space with explicit neighborhood structures. We posit that development along these directions will further improve the quality of co-construction and confirm HAI-Co2’s potential as an effective framework for the class of co-construction problems we aim to address. Modern software development relies on software engineering tools such as type systems, test drivers, static program analysis tools, monitoring and debugging tools and security vulnerability detectors. Realistic software artifacts are complex, and their full evaluation by humans without these tools is infeasible. Our implementation of HAI-Co2– not intended as a systematic evaluation of HAI-Co2’s effectiveness in the software domain – will need to be extended with many of the elements that are standard in the DevOps pipeline (see Le et al. (2022); Maninger et al. (2024) for examples of how to integrate such standard tools). We use the LLM’s generative capacity as is, e.g., when it explains why one generated solution is better than another. An interesting research direction would be explanatory interactive learning (Ross et al., 2017; Teso & Kersting, 2019; Friedrich et al., 2023), where more faithful explanations are produced through interactively constraining model explanations. The search strategy can be further refined by implementing reinforcement learning from execution feedback (Gehring et al., 2024; Liu et al., 2023; Dutta et al., 2024).
6 Conclusion
This paper presents a novel research perspective towards complex problem solving via human-AI co-construction. Our position is that existing generative AI agents require active human participation to successfully construct solutions to complex problems, but cannot effectively serve as reliable partners in human-AI cooperation due to their current limitations. We find evidence for this position in multiple prior research areas across a broad set of domains. Our case study focuses on software generation using GPT-4 Turbo, a strong proprietary LLM, and exemplifies the major drawbacks of current LLMs such as inability to follow human preferences, unreliable refinement of complex solution artifacts and limitations to facilitate active human participation. We observed that although day-to-day usage of generative AI tends to adopt a type of human-AI co-construction paradigm in an uninformed manner, the challenges that LLMs face confine such interactions to a much weaker form. As a remedy, we introduce HAI-Co2, a framework that is motivated by the effectiveness of collective human problem solving. HAI-Co2 facilitates a solution construction space with multiple levels of abstractions, in which human and AI iteratively refine the candidate solution through search guided by human preference. HAI-Co2 allows active human participation along with versatility in preference expression. After presenting a formalization of HAI-Co2, we discussed the research challenges for this new approach as well as possible future directions for addressing them. Finally, we presented a case study for the application of code generation and noted clear improvements compared to monolithic LLMs.
Appendix A Possible implementation in Related Work Generation

### v4

@@ 4 Challenges and future research
Given that human experts are costly and have limited time, LLM-based simulation of human-AI interaction may facilitate large-scale evaluation (Tamoyan, Schuff, and Gurevych, 2024). Even though our four evaluation criteria seem to demand human evaluation, we conjecture that the development of artificial critic models (McAleese et al., 2024), with human value alignment, will be an important research direction in the future.
5 Related work
We group relevant prior work into several major strands: human-AI cooperation, reinforcement learning from human feedback (RLHF), assistance games, learning from natural language interactions, search- and evolution-driven construction, inference-time compute scaling, LLM agents for complex problem solving, persistent solution spaces for iterative construction and human in the loop. We discuss how these approaches attempt to address (versions of) expert-AI co-construction. However, as we saw in Section 4, they fail to comprehensively tackle these challenges.
Human-AI Cooperation. Recent research emphasizes enhancing human-AI cooperation to support designers in complex, creative tasks. A notable example is the AI-assisted design (AIAD) framework proposed by De Peuter, Oulasvirta, and Kaski (2023), which mirrors our approach in emphasizing cooperation over automation and aiming to support designers’ creativity by inferring their goals. AIAD addresses the challenge of goal communication by using generative user models to understand designer reasoning and capabilities from their behavior. This allows AI to provide helpful recommendations and learn from the designer’s active participation and corrections. More broadly, this line of work aligns with the vision of Hybrid Intelligence (Akata et al., 2020), which advocates for human-AI systems that interact as co-evolving collaborators with complementary skills and adaptive behaviors. While both AIAD and Hybrid Intelligence share our goal of leveraging complementarity between human and machine, HAI-Co2 differs from them in several aspects. First, our approach explicitly adopts a hierarchical view encompassing multiple abstraction levels, like specifications, UML diagrams, and code, which is not a core element of AIAD or Hybrid Intelligence. Second, the search in our framework operates across this multi-level structure rather than within a monolithic or unstructured design space, enabling systematic refinement and revision. Third, we place stronger emphasis on actively learning not just from user feedback to AI suggestions, but also from direct human edits to evolving solution artifacts. Fourth, HAI-Co2 focuses on natural language communication for intuitive and flexible interaction between users and AI. Finally, whereas AIAD emphasizes minimally disruptive assistance with the aim of AI unobtrusively supporting designers, our framework envisions a more proactive and equal partnership, giving the AI system more responsibility in the co-construction process, including proposing refinements to the objective.
Reinforcement Learning from Human Feedback. RLHF focuses on learning a policy preferred by humans, most commonly relying on comparisons between candidate solutions (Kaufmann et al., 2023). The goal is to learn a policy that maximizes a reward or utility function that is consistent with the human feedback. Originating in classical reinforcement learning domains such as games and continuous control (Christiano et al., 2017), RLHF has been extended to a variety of domains, most notably fine-tuning generative models such as LLMs (Stiennon et al., 2020; Ouyang et al., 2022), eventually leading to the development of AI models such as ChatGPT that can generate human-preferred responses in natural language.
Assistance Games. Originally introduced as cooperative inverse reinforcement learning (Hadfield-Menell et al., 2016), assistance games (Shah et al., 2021; Laidlaw et al., 2024) model human-AI interaction as a game under partial information where the AI strives to learn and maximize the human’s underlying reward function. A key assumption is that humans have pre-defined objectives, even if initially unknown to the AI. While HAI-Co2 shares this overarching aim of AI-driven assistance for human experts in tackling complex problems via iterative engagement and the integration of human preferences, it fundamentally differs in its formalization and underlying assumptions.
An important distinction is the concept that the objective itself is not fixed in HAI-Co2 but co-constructed through interaction. In complex domains, expert preferences rarely remain static; they evolve through exploration and iterative refinement. This dynamic formation of objectives stands in contrast to the assistance game paradigm, where the AI primarily infers a static, pre-existing human reward function. Beyond this conceptual distinction, assistance games offer elegant theoretical properties but face practical challenges in real-world deployment due to computational complexity arising from the many possible combinations of AI and human policies as well as human objectives. [^2] Laidlaw et al. (2024) present steps toward scalably solving assistance games in more complex environments. While promising, our approach differs fundamentally through its emphasis on objective co-construction, structured search spaces, and natural-language communication. HAI-Co2 addresses these challenges by structuring the search process at multiple abstraction levels where humans provide feedback at their most intuitive level, while leveraging pretrained language models to guide the search process. While not suitable for all domains in which assistance games apply, we argue that this combination of structured solution spaces, natural language communication, and emphasis on objective co-construction provides significant advantages in the domain of expert-AI cooperation.
Learning from natural language interactions. Natural language interaction between the expert and the AI agent is central to HAI-Co2. A popular approach towards facilitating human-AI interactive problem solving involves training the agent to follow instructions in natural language (Branavan et al., 2009; Tellex et al., 2011). This paradigm of learning to follow instructions has found attention in the LLM era as well (Wei et al., 2022a). However, directly mapping language-specified goals to actions has limited applicability to novel tasks. Alternatively, learning rewards from natural language interaction to successfully align the AI agent with the human user has been explored (Fu et al., 2019; Sumers et al., 2021) – instead of learning to map language-defined goals to actions directly, they seek to learn the reward function from the language-defined goals that can be generalized to novel tasks. The findings of Sumers et al. (2022) suggest that while instructions typically perform well in low-autonomy settings, high-autonomy regimes favor the reward-learning paradigm. Instead of learning a language-conditioned policy or reward, it is also possible to use conversational cues as rewards themselves (Jaques et al., 2020), which can be combined with the other approaches discussed here. Yet another approach, deployed in OpenAI’s ChatGPT, is to enable the language model to save information about the user and their interactions with the model functioning as a natural-language ‘memory’, which may include information about the user’s preferences. [^3] https://openai.com/index/memory-and-new-controls-for-chatgpt/ Most of these approaches do not consider the need to actively elicit and co-construct the expert’s preferences, a key aspect of HAI-Co2. This necessity is supported by Co-Reyes et al. (2019) and Lin et al. (2022), who identify that inferring the correct behavior or reward from a single utterance is non-trivial given the multidimensionality of language. Querying the human user and estimating their preferences in an interactive setup is also a key component of Peng et al. (2024)’s framework. As a step in this direction, Li et al. (2023a) use active elicitation to strengthen preference understanding – the AI agent is trained to elicit and infer human preferences by actively interacting with the user. This is crucial prior work for an implementation of HAI-Co2, forming an important pillar of future research under the abstract umbrella it provides.
Search- and Evolution-Driven Construction. Our framework emphasizes iterative search within the construction space, a process akin to evolutionary optimization, which iteratively generates and evaluates candidate solutions (Bäck, 1996). This evolution can be viewed as a form of search-based construction. Interactive evolutionary computation, a preference-based extension, is particularly relevant to our work as it involves human evaluation of candidate solutions (Takagi, 2001; Wang and Pei, 2024). For example, these methods have been applied to search-based procedural content generation in video games (Togelius et al., 2011). Our approach differs in the core approach to the search process: Traditional evolutionary methods maintain a population of candidate solutions and generate new ones through mutation and recombination. In contrast, in our framework, each iteration ends with a single candidate solution that is then the basis for the next iteration. In addition, we leverage the extensive prior knowledge of pretrained language models to guide the search and use natural language communication to facilitate cooperation between the AI agent and the human expert.
Inference-time compute scaling. Allocating additional compute at inference-time can significantly enhance language models’ performance on complex tasks. Strategies vary along two axes: (1) depth vs. breadth (sequential refinement vs. independent exploration) and (2) structured vs. learned process control (externally imposed rules and algorithms vs. autonomous model capabilities guided by prompts or training). Depth-based methods encompass structured refinement, like systematic revision (Qu et al., 2024), and learned refinement, such as autonomous reasoning steps via chain-of-thought (Wei et al., 2022b; Kojima et al., 2022). Breadth-based methods often involve structured parallel generation with verification (e.g., best-of-N, Cobbe et al., 2021; Lightman et al., 2024). Recognizing complementary strengths of depth- and breadth-based approaches (Snell et al., 2024), recent work advocates combined approaches, either explicitly structuring integration (Wang et al., 2022; Yao et al., 2023; Snell et al., 2024) or relying on models’ learned or emergent self-reflection and adaptive search-like behaviors (OpenAI et al., 2024; DeepSeek-AI et al., 2025). Although both breadth-based parallel efforts and depth-based iterative revisions implicitly or explicitly search the solution space, this search is typically conducted non-interactively. This limits the ability to adapt generation to nuanced or shifting user goals during the construction process, relying solely on pre-defined, static objectives or verifiers, with minimal incorporation of explicit human preferences or feedback during search.
In contrast to these predominantly non-interactive methods optimized for static objectives, our approach introduces an interactive, preference-guided search paradigm explicitly designed to handle underspecified and evolving goals. While learned reasoning techniques like chain-of-thought excel in domains with clear objective correctness, such as mathematical or symbolic reasoning (Sprague et al., 2025), our method targets complex expert tasks where subjective preferences and evolving requirements critically shape optimal solutions. Instead of seeking solutions based on fixed objectives, we iteratively co-construct solutions aligned with dynamic expert preferences by explicitly integrating iterative feedback. This allows the search process to dynamically reshape the model’s reasoning trajectory in response to evolving goals, enabling alignment with expert decision-making in nuanced tasks. Beyond this distinction, many of the approaches discussed above are complementary to our framework and could help produce better base models or be used as inspiration for the search process itself.
LLM agents for complex problem solving. The rapid increase in the capabilities of LLMs has triggered multiple recent efforts to integrate them at the core of autonomous agents that interact with the environment, plan, and act to solve complex problems (Wang et al., 2024). Typical approaches adopt integrating different tool-usage capabilities into LLMs via efficient prompting, often with multimodal capabilities (Chen et al., 2023). A single agent is often insufficient to solve complex problems; thus, multiple agents with different capabilities have to be integrated. Recent efforts in LLM-based multi-agent systems seek to mimic such cooperative problem solving by role-playing LLMs via in-context examples (Li et al., 2023b) or fine-tuning (Juneja, Dutta, and Chakraborty, 2024). Multi-agent systems have started gaining traction in expert domain applications as well, e.g., software development (Qian et al., 2024), finance (Ganesh et al., 2024), chemistry (Song et al., 2025). The ChatDev workflow proposed by Qian et al. (2024) showcases the effectiveness of using natural language as the primary communication medium between roleplaying agents for software development. This demonstration in the context of AI-AI interaction strengthens our argument for natural language-based human-AI cooperation. These frameworks of LLM-based autonomous agents are largely designed toward AI autonomy. However, the target problems are not isolated applications – they require interactions between human organization(s), and as a result, these autonomous agents end up as incomplete assistants. Recently, arguments in favor of strategically allocating tasks between humans and LLM-based agents to exploit their distinct strengths have been put forward (He, Treude, and Lo, 2024), which aligns with our approach of leveraging the strengths of both humans and AI agents in co-construction. We argue that the open challenges in efficient interaction between a human and an agentic ecosystem, as outlined by Bansal et al. (2024), share some of the characteristics of human-AI co-construction as we envision it. An alternative view of our framework [^4] Indeed, our running example for HAI-Co2 is closely analogous to multi-agent systems. is hence an extension of LLM-based multi-agent systems with a human agent as a key component, focusing on the co-construction of solutions.
Human in the Loop (HIL). In its original form, HIL refers to a type of interactive machine learning system in which the human has the role of an annotator, advisor or provider of feedback that the system can solicit input from on specific questions and requests Mosqueira-Rey et al. (2023). For example, in active learning, the machine learning system identifies informative examples for the human to label and retrains the machine learning model iteratively until a termination criterion has been met. While there is no generally accepted definition of HIL and quite heterogeneous approaches have been grouped under this umbrella term by different authors McCarthy (1959); Settles (2009); Towell and Shavlik (1994); Schramowski et al. (2020); Dejong and Mooney (1986); Natarajan et al. (2014); Maclin et al. (2005); Boutilier (2002), HIL work includes some form of human feedback in contrast to completely automatic forms of machine learning.
In traditional HIL, the human’s role is limited: it is confined to the role of an oracle that is consulted with specific requests. The human is not involved in higher-level decisions or co-construction of a solution.
In contrast, as we discussed in Section 2, HAI-Co2 is a framework in which human and agent cooperate as equal partners, each fully engaged in co-constructing the solution to the problem and in co-constructing the objective. Thus, the cooperation between human and agent is more symmetric in HAI-Co2 than in HIL.
The need for a more equal partnership between human and AI agent – going beyond a narrow definition of HIL – is being recognized more generally. For example, Natarajan et al. (2025) argue that while there are problems in which the limited role of humans typical of traditional HIL is appropriate, many other problems (similar to the complex expert-domain problems we target) require a more equal partnership.
The term co-construction. We have borrowed the term co-construction from other disciplines, namely from the social sciences and humanities, where it usually refers to the “joint creation of form, interpretation, stance, action, activity, identity, [..] or other culturally meaningful reality” (Jacoby and Ochs, 1995). Although the definition allows for various possible interpretations, co-construction is interactional at its core. However, as the theory of co-construction is beyond the scope of this article and also out of our expertise, we would refer to Robertson et al. (2024, Section 3) who thoroughly discuss constructivist principles and human-AI knowledge co-construction from a theoretical perspective. Beyond this theoretical basis, our scope is substantially different from Robertson et al. (2024) who solely consider efficient techniques for hand-crafting prompts for business managers.
6 Conclusion
Our position is that existing generative AI agents require active human participation to successfully construct solutions to complex expert-domain problems, but cannot effectively serve as reliable partners in human-AI cooperation due to their current limitations. We find evidence for this position in prior research across a broad set of domains. Our running example focuses on software generation using GPT-4 Turbo, a strong proprietary LLM, and exemplifies the major drawbacks of current LLMs such as inability to follow human preferences, unreliable refinement of complex solution artifacts and limitations to facilitate active human participation. We observed that although day-to-day usage of generative AI tends to adopt a type of human-AI co-construction paradigm in an uninformed manner, the challenges that LLMs face confine such interactions to a much weaker form.
As a remedy, we introduce HAI-Co2, a framework that is motivated by the effectiveness of collective human problem solving. HAI-Co2 facilitates a solution construction space with multiple levels of abstractions, in which human and AI iteratively refine the candidate solution through search guided by human preference. HAI-Co2 allows active human participation along with versatility in preference expression. After presenting steps towards a formalization of HAI-Co2, we discussed the research challenges – including ethical challenges – for this new approach as well as possible future directions for addressing them.
Appendix A An exemplary simulation of HAI-Co2
@@
In this section, we present an example implementation of the major elements of HAI-Co2, tailored to code generation as a co-construction problem. This example does not claim scientific rigor on its own; instead, we use it to demonstrate what prior findings (see Section 1 and Section 5) already establish. We do not provide a complete implementation of HAI-Co2; in particular, the following are not included in the case study: neighborhood structure of the solution space, preference extraction from actual human participation, dedicated utility function tailored toward the expert problem. Instead, we emulate the intended behavior of a complete implementation using prompted LLMs, with the goal of motivating the practicality of HAI-Co2.
@@
To facilitate the exploration of the candidate solution space, we generate multiple solution representations on different abstraction levels by setting a high decoding temperature in the respective generation modules and sampling multiple responses. Intuitively, we seek to exploit earlier findings that a highly stochastic generation regime facilitates better novelty (Wang, Liu, and Awadallah, 2023). Furthermore, this imposed stochasticity can be interpreted as the notion of neighborhood in the respective spaces: one can treat two solutions sampled from the same input context of the generation module as neighbors; distance between two different input contexts can be measured by edit distance. Although we do not explicitly specify such geometric structure in this example, the search strategy uses it implicitly.
@@
The preference-based search among the candidate UMLs is implemented as a tournament by iteratively declaring one among a pair as the winner of a round. After a logarithmic order of such rounds ($\log_{2}n$ being the depth of the tournament tree for $n$ candidate UMLs), the Preference-based search module comes up with a final pair and a summary of preference justifications. [^6] See https://subha0009.github.io/ExAIC-Interactions/PreferenceLayer.html for the tournament on the candidate UMLs generated Note that although we seek to minimize the cost of human intervention in this step by automating preference-based ranking, one can envision the human expert providing their judgement on these UMLs. In such a setup, the Preference-decoding module can be used to explicitly adapt to such gold preference examples.
@@
We recruited a total of 14 participants for this evaluation; each is a doctoral student in Natural Language Processing, so that expert preferences in code generation can be understood. We provide each evaluator with a pair of interactions between i) a user and GPT-4 Turbo  [^7] Example interaction can be found at https://subha0009.github.io/ExAIC-Interactions/Assistant2.html and ii) a user and HAI-Co2  [^8] Example interaction can be found at https://subha0009.github.io/ExAIC-Interactions/Assistant1.html. Figure 4 shows the interface used to collect the users’ responses. We do not collect any personal information from the evaluators.
@@
On all five criteria, the majority of evaluators rate HAI-Co2 higher than GPT-4 Turbo. 12 (85.7%) evaluators find that HAI-Co2 better captures the initial preferences of the user (Q1). 11 (78.6%) agree that it can adapt better to preference switching (Q2). 11 (78.6%) see HAI-Co2 as superior on precision (Q3) and completeness of modification (Q4). 12 (85.7%) suggest HAI-Co2 is better suited for software development (Q5). Detailed responses are available here: https://subha0009.github.io/ExAIC-Interactions/FormResponses.html.
Figure 4: Interface used for human evaluation.
@@
Limitations and further improvements. The immediate improvements we observe are prevalent without any dedicated implementation – neither of the two refinement maps (from natural language to UML and from UML to Python) nor of the preference-based search policy. In this implementation, we do not equip the co-construction space with explicit neighborhood structures. We posit that development along these directions will further improve the quality of co-construction and confirm HAI-Co2’s potential as an effective framework for the class of co-construction problems we aim to address. Modern software development relies on software engineering tools such as type systems, test drivers, static program analysis tools, monitoring and debugging tools and security vulnerability detectors. Realistic software artifacts are complex, and their full evaluation by humans without these tools is infeasible. Our implementation of HAI-Co2– not intended as a systematic evaluation of HAI-Co2’s effectiveness in the software domain – will need to be extended with many of the elements that are standard in the DevOps pipeline (see Le et al. (2022); Maninger, Narasimhan, and Mezini (2024) for examples of how to integrate such standard tools). We use the LLM’s generative capacity as is, e.g., when it explains why one generated solution is better than another. An interesting research direction would be explanatory interactive learning (Ross, Hughes, and Doshi-Velez, 2017; Teso and Kersting, 2019; Friedrich et al., 2023), where more faithful explanations are produced through interactively constraining model explanations. The search strategy can be further refined by implementing reinforcement learning from execution feedback (Gehring et al., 2024; Liu et al., 2023; Dutta et al., 2024).
Appendix B Possible implementation in Related Work Generation